# Stochastic Two-temperature Nonequilibrium Ising model

### v1

@@ V Conclusion
We find that the NESS admits an effective temperature description for fast temperature switches, i.e., for large values of $\gamma$. We explicitly compute this effective temperature in the $\gamma\to\infty$ limit. Using extensive numerical simulation, we show that the effective temperature description remains valid for a finite range of $\gamma$ and breaks only in the small $\gamma$ regime. We further show that such an effective temperature description does not imply that the system reaches an equilibrium state—the inherent nonequilibrium nature of the dynamics leads to a non-zero energy current flowing through the system from the hot reservoir to the cold reservoir.
It would be interesting to investigate how the behaviour of the NESS changes as the system size is varied. In particular, how the threshold values of the flip rate and the effective temperatures depend on the system size. Another natural open question is whether similar effects are observed in other systems that belong to the Ising universality class. In a more general sense, our study can be thought of as a special case of a driving protocol where a control parameter of a system is randomly switched between two distinct values according to a dichotomous process. In particular it would be interesting to study the behaviour of systems with underlying nonequilibrium dynamics like the Katz-Lebowitz-Spohn model [32] or the driven lattice gas [35].
@@ Appendix A Response coefficient in the limiting case of infinite switching rate
In this appendix, we derive the expressions for second-order response coefficient $\chi_{2}(t)$ in the limiting scenario of infinite switching rate, i.e., for $\gamma\to\infty$. We start by considering the $\gamma$ dependent part of the response coefficient quoted in Eq. (26),
@@
in Eq. (26), where $\mathscr{I}(t)$ denotes the instantaneous energy current flowing into the system. In the unperturbed dynamics, i.e., for $\delta=0$, the dichotomous process $\sigma\to-\sigma$ and the Ising spin-flip dynamics at temperature $T_{c}$ become statistically independent, allowing the averages over $\sigma$ and over the spin configurations to be factorised. Consequently, Eq. (A.1) now simplifies to
@@
Substituting Eq. (A.4) in Eq. (A.3) and subsequently performing the integral over $t^{\prime\prime}$, yields

### v2

@@ 6 Conclusion
We find that the NESS admits an effective temperature description for fast temperature switches, i.e., for large values of $\gamma$. We explicitly compute this effective temperature in the $\gamma\to\infty$ limit, which turns out to be below the critical temperature of the ordinary Ising model. Using extensive numerical simulation, we show that the effective temperature description remains valid for a finite range of $\gamma$ and breaks only in the small $\gamma$ regime. We further show that such an effective temperature description does not imply that the system reaches an equilibrium state—the inherent nonequilibrium nature of the dynamics leads to a non-zero energy current flowing through the system from the hot reservoir to the cold reservoir. We also show that the qualitative behaviour of the observables in the NESS remains robust across system size variation.
A natural open question is whether similar effects are observed in other systems that belong to the Ising universality class. In a more general sense, our study can be thought of as a special case of a driving protocol where a control parameter of a system is randomly switched between two distinct values according to a dichotomous process. In this context, it would be worthwhile to investigate systems with underlying nonequilibrium dynamics like the Katz-Lebowitz-Spohn model [36] or the driven lattice gas [39]. It would also be of interest to explore whether our results can be tested experimentally using thin-film set-ups similar to the ones used in Ref. [40, 41].
@@ Appendix A Response coefficient in the limiting case of infinite switching rate
In this appendix, we derive the expressions for second-order response coefficient $\chi_{2}(t)$ in the limiting scenario of infinite switching rate, i.e., for $\gamma\to\infty$. We start by considering the $\gamma$ dependent part of the response coefficient quoted in (26),
@@
in (26), where $\mathscr{I}(t)$ denotes the instantaneous energy current flowing into the system. In the unperturbed dynamics, i.e., for $\delta=0$, the dichotomous process $\sigma\to-\sigma$ and the Ising spin-flip dynamics at temperature $T_{c}$ become statistically independent, allowing the averages over $\sigma$ and over the spin configurations to be factorised. Consequently, (A.1) now simplifies to
@@
Substituting (A.4) in (A.3) and subsequently performing the integral over $t^{\prime\prime}$, yields